How do you solve 3x^2+3=10x ?
Answer by Cromlix(4381)

Sort it out into ax^2 + bx + c = 0
3x^2 + 3 = 10x
3x^2 - 10x + 3 = 0
(3x - 1)(x - 3)= 0
3x - 1 = 0
x = 1/3
OR
x - 3 = 0
x = 3
Hope this helps.
:-)